An firm borrowed $164,000 to remodel their office. The load was to be paid back in equal monthly payments over 30 years at 5% interest, compounded monthly. After 13 years, the firm wants to pay off the loan. What is the remaining balance that must be paid off? Express your answer in $ to the nearest $1,000.
Answer #1

Initial loan = 164000

i = 5% compounded monthly = 5%/12 = 0.41667% monthly

t = 30 years = 30*12 = 360 months

Monthly payments = P*(A/P, 0.41667%, 360)

= 164000 * 0.005368219

= 880.388

Now after 13 years company wants to pay off the loan, we need to find principal left to be paid

Total payment periods left = (30-13) * 12 = 17*12 = 204 months

Principal left = A *(P/A, 0.41667%, 204)

= 880.388 * 137.239067

= 120823.61

= 121000 (rounding to nearest 1000)